two consectutive odd intergers have a sum of 152. what are the integrgers?
Answer by checkley71(8403)

X+X+2=152
2X+2=152
2X=152-2
2X=150
X=150/2
X=75 ANSWER FOR THE FIRST INTEGER.
75+2=77 FOR THE SECOND INTEGER.
PROOF
75+77=152
152=152